OpenStudy (anonymous):

What is 4z-3 when z = −3? ---- z+6

OpenStudy (mathmale):

You are given an algebraic fraction and are asked to evaluate it at a certain z value (z=-3). Show your work, please.

OpenStudy (sushi121212):

Just plug -3 in for z: 4(-3)-3 -12-3 = -15 -3+6=3

OpenStudy (anonymous):

@Sushi121212

OpenStudy (anonymous):

so the answer is 3? :/

OpenStudy (sushi121212):

For z+6, the answer will be 3 because you plug in -3: z+6= -3+6, which is like 6-3, so the answer is 3.

OpenStudy (anonymous):

umm i need the answer for the whole equation

OpenStudy (sushi121212):

What's the whole equation? Is 4z-3/z+6 a fraction?

OpenStudy (anonymous):

yes

OpenStudy (sushi121212):

Okay: 1. 4z-3 ____ (Plug in -3) z+6 2. 4(-3)-3 ______ (-3)+6 3. -12-3 ____ 3 4. -15 __ 3 5. -5 -5 is your answer.

OpenStudy (anonymous):

@Sushi121212

OpenStudy (anonymous):

are you sure it's -5 cuz i got 1.5

OpenStudy (sushi121212):

Could you show me what you did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (sushi121212):

For the 4*-3=1, you have to multiply 4 be -3 because there was a z there. 4 x -3 equals -12. Then, subtract -12 by -3 and you get -15 because if you have 2 negatives, you add them and add a negative sign. The -3+6 is correct, but at the end, you get: -15 ___ 3 The fraction sign is also a division sign. You divide -15 by 3 and get -5.